OpenStudy (anonymous):

can anyone break down this by steps? i need to find dy/dx by implicit differentiation

OpenStudy (anonymous):

\[2\sqrt{x}+\sqrt{y}=5\]

OpenStudy (anonymous):

rearrange this to 2(x)^(1/2) + y^(1/2) -5=0 differentiate each term to get dy/dx= x^(-1/2)+(1/2)y^(-1/2)